Solve for x
(-2 x/7) - 3 = -2(x 7)
a. -21/5
b. 5
c. -3
d. -5

Answer by chndnbl(31)

(-14 x)/7 - 3 = -2x - 14
(-14 x - 21) / 7 = -2x - 14
x - 35 = -14x - 98
x 14x = 35 - 98
15x = -63
x = -63 / 15
x = -21 / 5.